OpenStudy (anonymous):

Which is the 25th term of the arithmetic sequence where a1 = 8 and a9 = 64?

OpenStudy (anonymous):

\[a _{n}=a _{1}+(n-1)d\] this is general formula for any arithmetic seequence term. So using it for the 9th term we can find commun diference d: \[64=8+(9-1)d\] d=56/8=7 now the 25th term is: \[a _{25}=8+24*7=176\]

OpenStudy (anonymous):

What is the solution to the rational equation 4/3 plus the fraction 2 over x plus 5 = negative 6 over 3x plus 1

OpenStudy (anonymous):

not even thankyou for the previous one?

OpenStudy (anonymous):

it didn t send ? i apologize dearely my computer is stupid. thank you for your help i appreciate it

OpenStudy (anonymous):

(4/3)+2/(x+5)=-6/(3x+1)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

left side: \[4(x+5)+6\over 3(x+5)\]= \[-6\over(3x+1)\] now find comun denominator and add

OpenStudy (anonymous):

ok i got an answer. what did you get

OpenStudy (anonymous):

(4x+26)(3x+1)+(3x+15)6=0 12x\[12x ^{2}+4x+78x+26+18x+80=0\] \[12x ^{2}+100x+106=0\] solve this